Part III: A leading consumer packaged goods company is analyzing the impact of various advertising media on its sales. It hasa) Write the linear model equation for the regression line identified by this Regression Report.

b) Do all three advertising types (Television, Radio or Google) have a positive impact on sales? Which has the greatest impact? Explain your answer.

c) What is the confidence interval and p value for the independent variable “Radio”? Does radio have a significant linear relationship with Sales?

d) What is the net effect on Same-Week Sales of spending $1,000 in a week of television advertising? Show your work.

e) What would Same-Week Sales be if the company spent $10,000 on television ads, $8,000 on radio ads, and $4,000 on Google Ads? Show your work.
Answer #1

a) Regression equation is: sales= -9003.79 + 3.08 * television + 1.97 * radio + 4.08 * google

b) Yes, all three advertising types ( televison, radio, google) have a positive impact on sales, since, the slope of all the three advertising types are positive i.e slope of televison, radio, google are 3.08, 1.97, 4.08 respectively.

c) confidence interval for radio is (1.825, 2.124) and p value is zero. Yes, radio have a significant linear relationship with sales as the p value is zero and smaller the p-value higher is significance.

d) Since the slope of television is 3.08. Thus, unit increase in television ads result in additional $3.08. So, keeping the radio and google ads constant, and spending $1,000 in a week will result in net increase of $3080(3.08*1000) in sales.

e) sales = -9003.79 + 3.08 * (10,000) + 1.97 * (8,000) + 4.08 * (4000) = -9003.79+3080+15760+16320= 26156.21